OpenStudy (anonymous):

x-4/4 times 12x/x^2 -16

Parth (parthkohli):

\[\rm {x - 4 \over 4} \cdot {12x \over x^2 - 16} = {12x(x - 4) \over 4(x^2 - 16)} = {12x\cancel{(x - 4)} \over 4(x + 4)\cancel{(x - 4)}} = {\cancel4 \cdot 3x\over \cancel4(x + 4)}\]